OpenStudy (anonymous):

Using complete sentences, explain how to graph y = 1/2 x - 3 using the slope-intercept method

OpenStudy (anonymous):

hello?

OpenStudy (jhannybean):

start with your y-intercept -3, you know that this is an increasing function because the slope is positive.  your slope tells you which direction your graph is increasing and taking your slope, you can graph the line that way xD Another way is to start at your y-intercept, -3, and then look for your x-intercept and connect the two points, making sure your slope is increasing up by 1, and over by 2. in order to find our x-intercept, we make y=0. \[\large y= \frac{1}{2}x-3\]\[\large 0 = \frac12 x - 3\]\[\large 3 = \frac12 x\]\[\large 3(2) = \frac12 (2) x\]\[\large x=6\] Now plotting the x and y-intercepts, we can conclude that the line has a slope of 1/2, meaning from either intercept. (x or y) you can go across 2 units, and down 1 to end up at your next point. The graph will also pass through the points x=6 and y=-3

OpenStudy (jhannybean):



OpenStudy (anonymous):

thank you!

OpenStudy (precal):

up 1 and right 2 for the slope

OpenStudy (anonymous):

ok thanks